OpenStudy (anonymous):

f(x)=3x^2+2x-4 find the function? how will I solve it?

OpenStudy (anonymous):

split the middle term

OpenStudy (anonymous):

how to solve it? can you do the first step??

OpenStudy (anonymous):

do I have to put the value of x?

OpenStudy (anonymous):

how it is to prove its a function

OpenStudy (anonymous):

no srry it can be done by quadratic formula

OpenStudy (anonymous):

3x^2-2x-4=0 9x-2-2x-4=0 +2x +2x 11x-2-4=0 +2 +2 11x-4=0 +4 +4 11x=4 ------ ------ 11 11 x= 4 hope u got it

OpenStudy (anonymous):

but I have given the value of x

OpenStudy (anonymous):

so wat we have to find ??

OpenStudy (anonymous):

I have given x=0 . x=1 , x=-1 , x=-x ? I am so confused . Do I have to put all the values

OpenStudy (anonymous):

first of all u tell me u r in which class ?

OpenStudy (anonymous):

I am in algebra and trignomerty

OpenStudy (anonymous):

ok then u have to put all the values of x

OpenStudy (anonymous):

plz give medals

OpenStudy (anonymous):

thanxx.

OpenStudy (anonymous):

-f(x)

OpenStudy (anonymous):

so there are 6 values of x .so I have to put it six times?

OpenStudy (anonymous):

and solve it

OpenStudy (anonymous):

yess you have to do it its easy, it will take hardly 10 minutes

OpenStudy (anonymous):

what to do after finding the answerS? make a graph?

OpenStudy (anonymous):

is it given in the question to make a graph ???

OpenStudy (anonymous):

f(-1)=3x^2-2x-4

OpenStudy (anonymous):

yes u have to do it like this

OpenStudy (anonymous):

but put the values of x also in equation

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thats it

OpenStudy (phi):

f(x)=3x^2+2x-4 is shorthand f(x) is a number that depends on what x is. change x, and you get a different f(x) to find f(x) when x=0, f(0) you replace x in the definition and do the arithmetic f(0)= 3*0^2 + 2*0-4 (replace x with 0) 0^2 means 0*0 which is 0. 3*0 is 0. the 2nd term 2*0 is 0 so f(0)= 0+0-4= -4 (0,-4) is a point on this curve. (if you want to plot it) do the same for f(1):

OpenStudy (anonymous):

when x=1 then 5, -4 correct

OpenStudy (phi):

when x=1 then f(1)= 3*1*1+2*1-4 = 3+2-4 that simplifies to 5-4= 1 so f(1)=1 (1,1) is another pair

OpenStudy (phi):

try f(-1)

OpenStudy (anonymous):

when x is -x>

OpenStudy (anonymous):

sorry when x is -x?

OpenStudy (phi):

try f(-1)

OpenStudy (anonymous):

that is -1, -3

OpenStudy (anonymous):

when x is -x?

OpenStudy (anonymous):

3(-1)^2+2(-1)-4

OpenStudy (anonymous):

3-2-4

OpenStudy (phi):

and 3-2-4 is -3 Yes. so I take that back. You are correct. x = -x is only true for x=0 0=-0

OpenStudy (phi):

they want you to put -x into the equation f(x)= 3x^2 + 2x -4 so replace x with -x: f(-x)= 3*(-x)*(-x)+ 2*(-x) -4 f(-x)= 3x^2 -2x -4

OpenStudy (phi):

Any more ?

OpenStudy (anonymous):

after finding all the values do we have to make a graph or no?

OpenStudy (phi):

It depends on what the question is. Your function is a quadratic, representing a parabola.

OpenStudy (anonymous):

find function. Thats my question

OpenStudy (anonymous):

3x^2-2x-4 is the answer or do I have to solve more?

OpenStudy (phi):

f(x)= 3x^2-2x-4 is a function. Without more context, I can't give any more insight. If you want a plot, you can use the points you found. Check against wolfram: http://www.wolframalpha.com/input/?i=f%28x%29%3D3x%5E2%2B2x-4+

OpenStudy (anonymous):

so is this the answer?

OpenStudy (phi):

Does this question come with a chapter? Did they show an example? If someone asked me "find the function" and showed me f(x)= 3x^2-2x-4 I would say that is a parabola with a minimum value of -13/3 at x= -1/3 But if I saw what you were studying, I might answer differently.

OpenStudy (anonymous):

I am looking in the book and yes the question ends here in examples .

OpenStudy (phi):

can you figure out what they want?